[figure: 9.7 cm]

THE LOGARITHMIC RENORMALIZATION OF THE ONE-ELECTRON ENERGY IN HUBBARD MODEL

N.V. Ryzhanova^{a}, A.V. Vedyayev^{a}, M.Ye. Zhuravlev^{b} and P.B. Ziubin^{a} aMoscow State University, Phys. Depart., Vorob’jovy gory, Moscow 119899, Russia bInstitute of General & Inorganic Chemistry of RAS, 31 Leninskii prosp., Moscow,

117907, Russia

The self-energy of the one-electron Green function is investigated in the framework of diagrammatic technique for Hubbard operators. The electron-magnon interaction was taken into account in the diagrams of third order in electron hopping. The self-consistent equations for electron Green function were solved in logarithmic approximation. The logarithmic term in the imaginary part of one-electron self-energy was found. The problem of becoming heavier of electrons due to Hubbard repulsion is discussed.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : M. Ye. Zhuravlev

Address for correspondence : Institute of General & Inorganic Chemistry of RAS, 31 Leninskii prosp., Moscow, 117907, Russia

Email address : zhur@ionchran.rinet.ru